OpenStudy (anonymous):

If the area (in square units) of the region under the curve of the function f(x) = 5, on the interval from x = a to x = 8 is 20 square units, what is the value of a? 4 5 6 7 8

OpenStudy (anonymous):

a=4I If I express this problem in a figure,  \[(8 - a) \times 5 =20\] \[8 - a = 4\] a=4

OpenStudy (mathmale):

"If the area (in square units) of the region under the curve of the function f(x) = 5, on the interval from x = a to x = 8 is 20 square units, what is the value of a?" Another way to approach this problem is to recognize that it involves definite integration. Are you (David) familiar with that? The function being integrated is f(x)=5, and we want the area under this "curve" from x=a to x=8. The resulting area must be 20 square units. The necessary integral is \[\int\limits_{a}^{8}5dx,\] and integration results in